OpenStudy (anonymous):

What is the square of 4-7i?

zepdrix (zepdrix):

Hey Rose c: Welcome to OpenStudy. So when we square a binomial, we get the first term squared, twice product of both, and the second term squared.\[\Large\rm (4-7\mathcal i)^2\quad=\quad (4)^2+2(-7\mathcal i\cdot 4)+(-7\mathcal i)^2\]

zepdrix (zepdrix):

So we need to simplify a bit from there.

zepdrix (zepdrix):

What do you think Rosey? You can handle the first term I'm sure, how bout the other two?

OpenStudy (anonymous):

Yes but not the other two i need help with the whole problem

zepdrix (zepdrix):

The middle term simplifies down a bit if we just multiply the 2, 7 and 4.\[\Large\rm (4-7\mathcal i)^2\quad=\quad 16-2\cdot7\cdot4\cdot\mathcal i+(-7\mathcal i)^2\]Which gives us something like uhhhh,\[\Large\rm (4-7\mathcal i)^2\quad=\quad 16-56\mathcal i+(-7\mathcal i)^2\]ya? Then for the last term...

zepdrix (zepdrix):

\[\Large\rm (-7\mathcal i)^2=(-7)^2\mathcal i^2\]Remember that you're square the negative on the 7 as part of the multiplication, so that should be come positive. Remember what happens when you multiply i by itself?

OpenStudy (anonymous):

ok thanks